s 1) the function f(x)=x5-3x3/2+2 is given. Find the percent error by taking the amount of chang x=1, Ax=0,001
Answer #1


let y=f(s = x9_322 42 , x=1, B2=0.001 de = dwa 3 g 2 + 2 = 524_38) to even the sale - Zaxe =n20, constant =o dy = 564-92 cm= sy = 0.00050888519 sy = 0.0005089 Percentage Error = |dy, Gylx.oo AY - 10.0005 -0.0005089/ ... X 100 0.0005089 = 1-748870 %